One pipe can fill a tank in 40 hours, while a second pipe can fill the same tank in 32 hours. How long would it take both pipes working together to fill the tank?
Found 2 solutions by solver91311, richwmiller:
Answer by solver91311(24713)


If A can do a job in x time periods, then A can do of the job in 1 time period. Likewise, if B can do the same job in y time periods, then B can do of the job in 1 time period.

So, working together, they can do



of the job in 1 time period.

Therefore, they can do the whole job in:



time periods.



John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by richwmiller(17219)

x/40+x/32=1
x/5+x/4=8
4x+5x=160
9x=160
x=160/9